# DrVoice: Parallel Speech-Text Voice Conversation Model via Dual-Resolution Speech Representations

### v3

@@ 3.3 Training Strategy
CoM-Mixing Training. The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in Appendix B. During inference, specifying these system prompts enables the generation of desired output results. For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
@@
Core-Cocktail Training. We identify a learning rate dilemma: employing a high learning rate significantly compromises the performance of the MLLM, whereas using a low learning rate leads to training stagnation, with the loss decreasing quite slowly. To overcome this optimization challenge, we develop a specialized two-stage training strategy, termed “Core-Cocktail”. The first stage involves full fine-tuning the entire MLLM with a relatively high learning rate. The goal of this phase is to rapidly move the model’s parameters into a more favorable region of the loss landscape. However, to counteract the potential performance degradation of the MLLM caused by this aggressive initial training, an intermediate merging step is introduced. Drawing inspiration from Xiao et al. (2024), the parameters of the MLLM trained in the first stage ($M_{1}$) are merged with the parameters of the base, pre-trained LLM ($M_{0}$). This merging creates a new, interpolated model, $M_{r}$, as defined by the following equation:
@@ 4 Experiments
Table 2: Performance Comparison on various benchmarks in terms of benchmark-specific metrics (the best result in each row is in bold). With the exception of GLM4-Voice, whose results are cited from Xu et al. (2025) and Chen et al. (2024b), all other results were generated by running inference on the released checkpoints. FR(In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone. $\tau$ denotes the average number of text tokens corresponding to one second of speech.
GLM4-Voice
MiniCPM
-o 2.6
Baichuan
-Omni-1.5
Qwen2.5
-Omni
Kimi
-Audio
Step-Audio2
-Mini
DrVoice
@@
AlpacaEval 57.89 64.10 77.90 72.76 75.73 59.60 78.34
Llama Q. 76.00 78.00 78.50 75.33 79.33 75.00 80.33
Reasoning QA 47.43 38.60 50.00 63.76 58.02 46.04 57.92
TriviaQA 51.80 63.00 57.20 57.06 62.10 57.70 61.50
Web Q. 55.40 69.20 59.10 62.80 70.20 65.10 68.10
Overall 57.70 62.58 64.54 66.34 69.08 60.69 69.24
@@
AlpacaEval 3.97 4.42 4.50 4.33 4.46 4.17 4.52
CommonEval 3.42 4.15 4.05 3.84 3.97 3.00 3.77
SD-QA 36.98 50.72 43.40 57.41 63.12 56.06 68.54
MMSU 39.75 54.78 57.25 56.38 62.17 52.18 60.31
OpenBookQA 53.41 78.02 74.51 79.12 83.52 64.18 79.56
IFEval 52.80 49.25 54.54 53.88 61.10 38.01 59.30
AdvBench 88.08 97.69 97.31 99.62 100.00 93.08 98.65
Overall 59.83 71.69 71.14 72.83 76.93 63.84 76.02
@@
AlpacaEval 51.00 51.00 58.69 56.10 44.20 51.72 49.65
Llama Q. 50.00 61.00 67.33 66.30 57.33 67.67 68.00
TriviaQA 36.40 40.20 30.57 40.52 35.71 33.50 35.35
Web Q. 32.00 40.00 38.09 38.93 33.90 34.65 37.65
Overall 42.35 48.05 48.67 50.46 42.79 46.89 47.66
@@
Kimi-Audio (2025) 12.5/12.5 3.06 21.06
@@
DrVoice 5/5 4.29 11.2
Table 4: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing Training on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
@@ 4.1 Experimental Setup
Datasets. Approximately 100K hours of audio-text paired data for speech-text modality alignment is adopted for pre-training Speech Refined Head. For DrVoice post-training, we first synthesize speech for about 3B text tokens using CosyVoice (Du et al., 2024a, b, 2025), then select about 26K hours for speech-to-speech conversation and about 20K hours user speech plus 1.3B assistant tokens for speech-to-text conversation, based on Word Error Rate (WER) of the synthesized speech. Furthermore, to enhance the model’s comprehension of real-world speech, the training data was augmented with about 10K hours of English Automatic Speech Recognition (ASR) data, comprising several corpora, including Common Voice (Ardila et al., 2020), MELD (Poria et al., 2019), LibriSpeech (Panayotov et al., 2015), SPGISpeech (O’Neill et al., 2021), and Voxpopuli (Wang et al., 2021). Following prior works (Yao et al., 2024; KimiTeam et al., 2025; OpenAI, 2024b), we evaluate the performance on the widely used benchmarks, VoiceBench (Chen et al., 2024b) and OpenAudioBench  [^1] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation, UltraEval-Audio  [^2] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation, and Big Bench Audio  [^3] https://huggingface.co/datasets/ArtificialAnalysis/big_bench_audio for both evaluations.
Evaluation Metrics. Evaluations adhere to the established protocols for each respective benchmark. Specifically, for the open-ended QA tasks on AlpacaEval and CommonEval, G-Eval (Liu et al., 2023) is used for scoring. For AdvBench, the Refusal Rate is reported, while performance on all other benchmarks is assessed with Accuracy. The generated speech is transcribed using Whisper-v3-large model (Radford et al., 2022), then WER (denoted by ASR-WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text. UTMOS (Saeki et al., 2022) is used to evaluate the overall speech quality, following Zeng et al. (2025).
Baselines. Representive and competitive open-source audio language models are selected as baselines to cover diverse modeling paradigms: Text-Driven models including MiniCPM-o 2.6 (8B) (Yao et al., 2024) and Qwen2.5-Omni (7B) (Xu et al., 2025). Among Joint Speech-Text models, interleaved models including GLM-4-Voice (9B) (Zeng et al., 2025), Baichuan-Omni-1.5 (7B) (Li et al., 2025), and Step-Audio2-Mini (8B) (Wu et al., 2025), alongside the parallel model Kimi-Audio (7B) (KimiTeam et al., 2025). This suite enables systematic comparisons across mainstream speech-text modeling strategies.
@@ 4.2 Main Results
Computational Efficiency and Speech Quality. A key innovation of DrVoice is its remarkable computational efficiency, highlighted in Table 2 and Table 3. As shown in the FR (In/Out) rows, DrVoice operates at a frame rate of 5/5, indicating that the LLM backbone processes only 5 audio tokens per second for both input and output. This is a substantial reduction compared to other models, which operate at frame rates of 12.5 or 25, thereby significantly lowering computational requirements and potential latency. Crucially, this efficiency does not come at the cost of speech quality. Table 3 shows that despite its low frame rate, DrVoice produces high-quality and well-aligned speech. It achieves UTMOS score of 4.29 for overall speech quality, which is competitive with top-performing models like Qwen2.5-Omni (4.28) and superior to others like Kimi-Audio (3.06). With an ASR-WER of 11.2, the model shows robust speech-text alignment and surpasses several baselines. Still, it underperforms compared to Qwen2.5-Omni. A potential explanation lies in the architectural design: Qwen2.5-Omni’s feeding text directly to its “Talker” module, whereas the proposed model only sends hidden states to SRH. Therefore, introducing text to the speech refined head is a logical next step for improving textual control and lowering the ASR-WER. This unique combination of SOTA performance, high-quality speech generation, and unparalleled efficiency makes DrVoice a highly powerful and practical model for real-world applications.
@@ 4.3 Ablation and Analyses
Continuous Speech Encoder. As shown in Table 4, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs. Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation. Specifically, the S2T score decreases from 72.33 to 62.33 (13.8% relatively), and the S2M (T) score falls from 68.67 to 61.67 (10.2% relatively). In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00. This confirms the CSE’s effectiveness in representing speech.
Dual-Resolution Speech Representations (DRSR). Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation. 1) Grouping Factor. Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M). For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain). The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5. Furthermore, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, proving the efficiency of grouping machinism. Detail results can be found in Appendix C. 2) Speech Refinement Head. As shown in Table 4, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks. By comparing the model with SRH (w/o. SRH-Pretraining) to one without (w/o. SRH), we observe that adding SRH achieves a 76.9% relative improvement in S2M (T) (from 21.67 to 38.33) and a 31.2% relative improvement in T2M (T) (from 45.22 to 59.33). The text-only performance (T2T) remains stable, confirming that SRH enhances speech generation without interfering with text processing pathways. Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
@@ 5 Conclusions
We introduce DrVoice, a novel parallel speech-text voice conversation model that leverages joint autoregressive modeling with dual-resolution speech representations. Our experimental results demonstrated that DrVoice establishes a new state-of-the-art (SOTA) on OpenAudioBench for audio understanding and on Big Bench Audio for reasoning capabilities. Furthermore, it achieved performance comparable to the SOTA on VoiceBench for voice assistant tasks and on UltraEval-Audio for both speech understanding and generation, confirming its strong and versatile capabilities. Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency without sacrificing performance. We discuss limitations and future work in Appendix D.
@@ Appendix A Implementation Details
Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT). The maximum sequence length is set to 2K tokens, which corresponds to an audio duration of approximately 6.8 minutes. DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. Core-cocktail interpolated factor is set to $\alpha=0$ for extremely maintaining the base LLM capability. Multiple loss hyperparameters are set to $\lambda=1$ and $\mu=1$. The warmup rate is set to $2\%$ of the total training steps. For the two-stage training of DrVoice, the learning rate is decayed from $1\times 10^{-4}$ to $1\times 10^{-5}$ in stage one, and subsequently from $2\times 10^{-5}$ to $2\times 10^{-6}$ in stage two, with both stages utilizing a cosine annealing schedule. In contrast, DrVoice-Small and SRH-PT are trained in a single stage, which adopts the same learning rate schedule as the first stage of DrVoice. The AdamW method (Loshchilov and Hutter, 2019) is used for optimization. Experiments are run on 64x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding. For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post-training.

### v4

@@ 3.3 Training Strategy
CoM-Mixing Training. The chain-of-modality (CoM) methodology (DBLP:conf/emnlp/ZhangLZZWZQ23) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in Appendix B. During inference, specifying these system prompts enables the generation of desired output results. For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
@@
Core-Cocktail Training. We identify a learning rate dilemma: employing a high learning rate significantly compromises the performance of the MLLM, whereas using a low learning rate leads to training stagnation, with the loss decreasing quite slowly. To overcome this optimization challenge, we develop a specialized two-stage training strategy, termed “Core-Cocktail”. The first stage involves full fine-tuning the entire MLLM with a relatively high learning rate. The goal of this phase is to rapidly move the model’s parameters into a more favorable region of the loss landscape. However, to counteract the potential performance degradation of the MLLM caused by this aggressive initial training, an intermediate merging step is introduced. Drawing inspiration from DBLP:conf/acl/XiaoLZX24, the parameters of the MLLM trained in the first stage ($M_{1}$) are merged with the parameters of the base, pre-trained LLM ($M_{0}$). This merging creates a new, interpolated model, $M_{r}$, as defined by the following equation:
@@ 4 Experiments
Table 2: Performance Comparison on various benchmarks in terms of benchmark-specific metrics (the best and second-best results in each row are in bold and underlined, respectively). With the exception of GLM4-Voice, whose results are cited from DBLP:journals/corr/abs-2503-20215 and DBLP:journals/corr/abs-2410-17196, and the BBA results, which are cited from  coreteam2025mimoaudio except for MiniCPM-o 2.6 that we reproduce, all other results were generated by running inference on the released checkpoints. FR(In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone. $\tau$ denotes the average number of text tokens corresponding to one second of speech.
GLM4-Voice MiniCPM -o 2.6 Baichuan -Omni-1.5 Qwen2.5 -Omni Kimi -Audio Step-Audio2 -Mini DrVoice
@@
AlpacaEval 57.89 64.10 77.90 72.76 75.73 59.60 82.61
Llama Q. 76.00 78.00 78.50 75.33 79.33 75.00 83.00
Reasoning QA 47.43 38.60 50.00 63.76 58.02 46.04 59.90
TriviaQA 51.80 63.00 57.20 57.06 62.10 57.70 64.50
Web Q. 55.40 69.20 59.10 62.80 70.20 65.10 70.20
Overall 57.70 62.58 64.54 66.34 69.08 60.69 72.04
@@
AlpacaEval 3.97 4.42 4.50 4.33 4.46 4.17 4.74
CommonEval 3.42 4.15 4.05 3.84 3.97 3.00 4.08
SD-QA 36.98 50.72 43.40 57.41 63.12 56.06 64.30
MMSU 39.75 54.78 57.25 56.38 62.17 52.18 67.27
OpenBookQA 53.41 78.02 74.51 79.12 83.52 64.18 82.20
IFEval 52.80 49.25 54.54 53.88 61.10 38.01 71.39
AdvBench 88.08 97.69 97.31 99.62 100.00 93.08 99.62
Overall 59.83 71.69 71.14 72.83 76.93 63.84 80.17
@@
AlpacaEval 51.00 51.00 58.69 56.10 44.20 51.72 59.29
Llama Q. 50.00 61.00 67.33 66.30 57.33 67.67 75.33
TriviaQA 36.40 40.20 30.57 40.52 35.71 33.50 46.09
Web Q. 32.00 40.00 38.09 38.93 33.90 34.65 45.92
Overall 42.35 48.05 48.67 50.46 42.79 46.89 56.66
@@
Kimi-Audio (kimiteam2025kimiaudiotechnicalreport) 12.5/12.5 3.06 21.06
@@
DrVoice 5/5 4.29 8.36
Table 4: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing on Llama Questions. S2M, S2T, etc are defined in Table 1.
@@ 4.1 Experimental Setup
Datasets. Approximately 100K hours of audio-text paired data for speech-text modality alignment is adopted for pre-training Speech Refined Head. For DrVoice post-training, we first synthesize speech for about 3B text tokens using CosyVoice (DBLP:journals/corr/abs-2407-05407; DBLP:journals/corr/abs-2412-10117; DBLP:journals/corr/abs-2505-17589), then select about 26K hours for speech-to-speech conversation and about 20K hours user speech plus 1.3B assistant tokens for speech-to-text conversation, based on Word Error Rate (WER) of the synthesized speech. Furthermore, to enhance the model’s comprehension of real-world speech, the training data was augmented with about 10K hours of English Automatic Speech Recognition (ASR) data, comprising several corpora, including Common Voice (DBLP:conf/lrec/ArdilaBDKMHMSTW20), MELD (DBLP:conf/acl/PoriaHMNCM19), LibriSpeech (DBLP:conf/icassp/PanayotovCPK15), SPGISpeech (DBLP:conf/interspeech/ONeillLMNZKBDFS21), and Voxpopuli (DBLP:conf/acl/WangRLWTHWPD20). Following prior works (yao2024minicpm; kimiteam2025kimiaudiotechnicalreport; openai2024hello), we evaluate the performance on the widely used benchmarks, VoiceBench (DBLP:journals/corr/abs-2410-17196) and OpenAudioBench  [^1] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation, UltraEval-Audio  [^2] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation, and Big Bench Audio  [^3] https://huggingface.co/datasets/ArtificialAnalysis/big_bench_audio for both evaluations.
Evaluation Metrics. Evaluations adhere to the established protocols for each respective benchmark. Specifically, for the open-ended QA tasks on AlpacaEval and CommonEval, G-Eval (DBLP:conf/emnlp/LiuIXWXZ23) is used for scoring. For AdvBench, the Refusal Rate is reported, while performance on all other benchmarks is assessed with Accuracy. The generated speech is transcribed using Whisper-v3-large model (radford2022whisper), then WER (denoted by ASR-WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text. UTMOS (DBLP:conf/interspeech/SaekiXNKTS22) is used to evaluate the overall speech quality, following DBLP:conf/iclr/ZengDLZJD025.
Baselines. Representive and competitive open-source audio language models are selected as baselines to cover diverse modeling paradigms: Text-Driven models including MiniCPM-o 2.6 (8B) (yao2024minicpm) and Qwen2.5-Omni (7B) (DBLP:journals/corr/abs-2503-20215). Among Joint Speech-Text models, interleaved models including GLM-4-Voice (9B) (DBLP:conf/iclr/ZengDLZJD025), Baichuan-Omni-1.5 (7B) (li2025baichuan), and Step-Audio2-Mini (8B) (DBLP:journals/corr/abs-2507-16632), alongside the parallel model Kimi-Audio (7B) (kimiteam2025kimiaudiotechnicalreport). This suite enables systematic comparisons across mainstream speech-text modeling strategies.
@@ 4.2 Main Results
Computational Efficiency and Speech Quality. A key innovation of DrVoice is its remarkable computational efficiency, highlighted in Table 2 and Table 3. As shown in the FR (In/Out) rows, DrVoice operates at a frame rate of 5/5, indicating that the LLM backbone processes only 5 audio tokens per second for both input and output. This is a substantial reduction compared to other models, which operate at frame rates of 12.5 or 25, thereby significantly lowering computational requirements and potential latency. Crucially, this efficiency does not come at the cost of speech quality. Table 3 shows that despite its low frame rate, DrVoice produces high-quality and well-aligned speech. It achieves UTMOS score of 4.29 for overall speech quality, which is competitive with top-performing models like Qwen2.5-Omni (4.28) and superior to others like Kimi-Audio (3.06). With an ASR-WER of 8.36, the model shows robust speech-text alignment and surpasses several baselines. Still, it underperforms compared to Qwen2.5-Omni. A potential explanation lies in the architectural design: Qwen2.5-Omni’s feeding text directly to its “Talker” module, whereas the proposed model only sends hidden states to SRH. This unique combination of SOTA performance, high-quality speech generation, and unparalleled efficiency makes DrVoice a highly powerful and practical model for real-world applications.
@@ 4.3 Ablation and Analyses
Continuous Speech Encoder. As shown in Table 4, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs. In this setting, user audio input is processed solely using the discrete speech tokenizer (same as the assistant’s output tokenizer), without the Whisper features. Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation. Specifically, the S2T score decreases from 72.33 to 62.33 (13.8% relatively), and the S2M (T) score falls from 68.67 to 61.67 (10.2% relatively). In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00. This confirms the CSE’s effectiveness in representing speech.
Dual-Resolution Speech Representations (DRSR). Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation. 1) Grouping Factor. Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M). For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain). The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5. Furthermore, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, proving the efficiency of grouping machinism. Detail results can be found in Appendix C. 2) Speech Refinement Head. As shown in Table 4, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks. To handle the resolution mismatch when SRH is removed, the hidden state from the shared LLM is passed through a projection layer ($d_{model}\to k\times d_{speech\_token}$) and then split into k tokens. This setup simulates a standard parallel prediction approach without autoregressive refinement. By comparing the model with SRH (w/o. SRH-Pretraining) to one without (w/o. SRH), we observe that adding SRH achieves a 76.9% relative improvement in S2M (T) (from 21.67 to 38.33) and a 31.2% relative improvement in T2M (T) (from 45.22 to 59.33). The text-only performance (T2T) remains stable, confirming that SRH enhances speech generation without interfering with text processing pathways. Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
@@ 5 Conclusions
We introduce DrVoice, a novel parallel speech-text voice conversation model that leverages joint autoregressive modeling with dual-resolution speech representations. Our experimental results demonstrated that DrVoice establishes a new state-of-the-art (SOTA) on OpenAudioBench for audio understanding, VoiceBench for voice assistant tasks, UltraEval-Audio for both speech understanding and generation and on Big Bench Audio for reasoning capabilities, confirming its strong and versatile capabilities. Notably, its dual-resolution mechanism significantly improves inference speed and computational efficiency (empirically achieving a reduction of nearly 50% in GPU hours for training) without sacrificing performance. We discuss limitations and future work in Appendix D.
[…]
@@ Appendix A Implementation Details
Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT). The maximum sequence length is set to 2K tokens, which corresponds to an audio duration of approximately 6.8 minutes. DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. Core-cocktail interpolated factor is set to $\alpha=0$ for extremely maintaining the base LLM capability. Multiple loss hyperparameters are set to $\lambda=1$ and $\mu=1$. The warmup rate is set to $2\%$ of the total training steps. For the two-stage training of DrVoice, the learning rate is decayed from $1\times 10^{-4}$ to $1\times 10^{-5}$ in stage one, and subsequently from $2\times 10^{-5}$ to $2\times 10^{-6}$ in stage two, with both stages utilizing a cosine annealing schedule. In contrast, DrVoice-Small and SRH-PT are trained in a single stage, which adopts the same learning rate schedule as the first stage of DrVoice. The AdamW method (DBLP:conf/iclr/LoshchilovH19) is used for optimization. Experiments are run on 64x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(DBLP:journals/corr/abs-1905-12322) to accelerate training and decoding. For DrVoice, DeepSpeed ZeRO-2 (DBLP:conf/sc/RajbhandariRRH20) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post-training.